OpenStudy (anonymous):

If the radius and height of a cylinder are multiplied by 4, the lateral area of the cylinder is multiplied by 2. 4. 8. 16.

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

Lateral Surface Area of a cylinder is \[2\pi\times r \times h\]